A monopoly faces a market demand Q(p)=1500-5p and has costs C(q)=120q. What is profit of the...

  1. A monopoly faces a market demand Q(p)=1500-5p and has costs C(q)=120q. What is profit of the firm?

    40500

    20250

    36000

    18000

    0
Answer #1

Now 300 - 2q = 120 1500-2q = 600 29 = goo [q=450] = profit - TR - TC - 300 q- q - 1209 a When q = 450 2 300x450- (450)2 - 120